Two long straight parallel wires, standing on air 2m apart carry currents i1 and i2 in the same direction. The magnetic intensity at a point mid way between the wires of 7.95 AT/m. If the force on each wire per unit length is 2.4 × 10-10. Evaluate i1 and i2
1
Expert's answer
2021-09-10T00:03:37-0400

Given:

"d=2\;\rm m"

"H=7.95\:\rm A/m"

"F/l=2.4*10^{-10}\:\rm N"


The magnetic intensity at a point mid way between the wires

"H=H_1+H_2=\frac{I_1}{2\pi d/2}+\frac{I_1}{2\pi d/2}"

"H=\frac{I_1+I_2}{\pi d}"

The force on each wire per unit length

"F/l=\mu_0\frac{I_1I_2}{2\pi d}"

Using these equations, we get

"I_1+I_2=2\pi*7.95\\
I_1*I_2=2*2.4*10^{-10}/10^{-7}"

"I_1+I_2=50\\
I_1*I_2=0.0048"

Roots:

"I_1=50\:{\rm A},\quad I_2=9.6*10^{-5}\:\rm A"